Sweeney originally contributed $310,000 in cash for a one-fourth interest in the Gilbert LLC. During the several years that Sweeney was a member of the LLC, his share of the LLC's income was $186,000 and he withdrew $124,000 cash. The LLC's liabilities are $155,000, of which Sweeney's share is $38,750. The LLC has $77,500 of hot assets, of which Sweeney's share is $19,375. Sweeney sells his LLC interest to Jana for $409,200 cash, with Jana also assuming Sweeney's share of the LLC's liabilities.

  1. How much is Sweeney's gain on the sale, and what is its character?

    Sweeney's total gain on the sale is $ , of which $ recognized as ordinary income, with the remaining $ recognized as a long-term capital gain.

    How much is Jana's adjusted basis for her LLC interest?
    $
Answer #1

ANSWER

409,200 +38,750=447,950

310,000 +186,000 -124,000 =372,000

372,000+38,750=410,750

Gain on sale=447,950-410,750=37,200

Adjusted basis=409,200 +38,750=447,950

_____________________________________________

If you have any query or any Explanation please ask me in the comment box, i am here to helps you.please give me positive rating.

*****************THANK YOU**************
Answer #2

How much is Ordinary income and how much is long term capital gain?